# Redefining Neural Operators in $d+1$ Dimensions

### v2

@@ 3.1 Model Architecture and Implementation
An overview of the SKNO architecture is shown in Fig. 3. Except for the lifting and measuring modules, we implement the SKNO by constructing signal propagators with the kernel integration in Equ.8, and with the aid of residuals. Each of them propagates the $d+1$ dimensional signal $\bm{v}_{l}$ with different inductive biases on global or local regions on $D_{x}$.
@@
In the Schrödingerisation process (Jin et al. (2023)) like Example 1 and 2, different differential operators with respect to $x$ (e.g. $\partial_{x}$, $\partial_{xx}$, …) can be represented by a complex diagonal matrix in Fourier domain of $x$ according to the properties of Fourier Transform.
Based on that, we leverage the implementation of spectral convolution operator in (Li et al. (2021)), which uses a learnable complex vector equivalent to a diagonal matrix in the Fourier domain of $D_{x}$, to parameterize the evolving operator related to $d+1$ dimensional PDEs. A typical example is shown in Remark 1. In practice, the complex vector is truncated for working on different grid sizes:
$SpectralConvOp(\bm{v}_{l})=F_{x}^{-1}(trun\_diag(F_{x}(\bm{v}_{l}))),$ (11)
where $F_{x}$ and $F_{x}^{-1}$ denote DFT and Inverse DFT (accelerated by Fast Fourier Transform (FFT) in practice). The trun_diag()trun\_diag() represents the element-wise multiplication between a learnable truncated complex vector and signal slices of $\bm{v}_{l}$ on the transformed grid $\xi$.
@@
Since the truncation in Sec.3.1 prefer to propagate signals in a relatively global scope, we compensate for possible local information lost by propagating the finite difference derivative of $x$ on the local region:
where $D_{loc}\subset D_{x}$ denotes the local region around the query position $x$. In this work, the implementation of Equ.12 is based on the convolution layer in (Liu-Schiaffini et al. (2024)).
@@
With above two implementations of $d+1$ dimensional kernel integral operators in Equ.8, we specified the linear block $\mathcal{L}$ in Def.1 with two parts: kernel integral operator $\mathcal{K}_{l}$ and residual operator $\bar{\mathcal{W}}_{l}$:
@@
Figure 4: Evolution along the auxiliary dimension and neural operator implementations. Left: The target operator $\mathcal{L}^{\prime}$ governs the evolution along pp. Right: Compared with other methods using discretized matrices for each position of $p$, our model captures the target evolution with operator modules in both position and momentum spaces.
@@ 3.2 Evolution along the Auxiliary Dimension
where $x_{{i}}=(x_{i_{1}},...,x_{i_{d}}),\ {i}=(i_{1},...,i_{d})$. Note that the same conclusion also works for $\xi_{i}$. And $K$ represents the highest order of the differential operator w.r.t $p$.
The evolution governed by Equ.14 is implemented by the forward finite difference method on $t=l$ (Let $\Delta t=1$ to simplify the equation, which can be adjusted during learning.) as follows:
@@
which is exactly our linear block in Equ.13. For discrete signal along $p$, we implement the right hand side of Equ.15 by ${F}_{p}^{-1}\tilde{A}{F}_{p}+\bm{b}$, where $\tilde{A}$ denotes a learnable matrix in Fourier domain on $p$ (convolution theorem) and $\bm{b}$ corresponds to the discretized $b(p)$. As illustrated in Fig.4, while other methods should adjust their discretization size of $p$ to capture the evolving dynamics in Equ.14, our model integrates both weighted differential operator represented in Fourier domain on $p$ (SKNO, right) and bias term (SKNO, left) to align itself with general linear PDEs as Equ.14. By iteratively updating the $d+1$ dimensional signal using the combination of these operator designs, our SKNO model aligns itself well with $d+1$ dimensional PDEs to effectively offer improved scalability and adaptability over previous methods.
@@ 4.1 General Setting
We evaluate our model on ten diverse benchmarks, covering a wide spectrum of PDE dynamics. To highlight the effectiveness of our evolving operator along the auxiliary dimension, we first test on simple linear dynamics, including the 1D Heat Equation (Shin et al. (2022)) and Advection Equation (Takamoto et al. (2022)). We then move to more challenging nonlinear and unstable regimes, conducting extensive experiments on the 1D Burgers Equation, 2D Darcy Flow (discontinuous coefficients), 2D incompressible Navier–Stokes (low viscosity 1e51e{-5}) (Li et al. (2021); Wu et al. (2024)), 2D Gray–Scott Reaction–Diffusion (Kassaï Koupaï et al. (2024)), and the highly unstable 3D Rayleigh–Taylor Instability (Ohana et al. (2024)). Additional benchmarks and detailed results are provided in Appendix C.
@@
We compare our model with various $d$-dimensional neural implementations, including DeepONet (Lu et al. (2021)), Fourier Neural Operator (FNO) (Li et al. (2021)), Convolutional Neural Operator (CNO) (Raonic et al. (2023a)), and transformer-based model Transolver (Wu et al. (2024)). All models are trained under their officially recommended optimal configurations, following the settings reported in their respective papers. For benchmarks such as 1D Burgers, 2D Darcy Flow, and 2D Gray–Scott, we train each model for 500 epochs with no physical prior regularization, using only relative $L_{2}$ loss. For the incompressible 2D Navier–Stokes, we ensure fairness by matching comparable training time scales, accounting for differences in training speed between transformer-based and other operator architectures. All experiments are run on a single Nvidia GeForce RTX 4090 GPU. Complete configurations are provided in Tab.15.
4.2 Main Results and Analysis
Results
Our SKNO model consistently outperforms existing methods, achieving the lowest relative $L_{2}$ error with competitive training costs. Tab. 6 summarizes the performance of different linear blocks $\mathcal{L}$ across representative models: the Fourier Integral Operator in FNO, Physics-Attention in Transolver, and our redefined Spectral Convolution Operator. All modules are evaluated with their original lifting and recovering components, without additional positional encoding (see Appendix C). For the heat equation experiment, we predict the solution at $t=1$ following Example 1, keeping the same setting as in practice with their mode or slice truncated. Note that these two parameterize the invertible aggregation on $D_{x}$. For the advection equation, we test these blocks in an autoregressive way to validate our proposed module’s performance on concatenated input quantities and prediction tasks. Specifically, we calculate the prediction error here of the autoregressive prediction in the next 10 time steps from the last 10. While other modules are limited by their evolving design on $p$, ours breaks this bottleneck and stably well captures the $d+1$ dimensional evolution. Across the benchmarks in Tab. 7, Tab. 18 and Tab. 20, SKNO achieves state-of-the-art (SOTA) performance, with training time comparable to FNO which is well-known for balancing speed and accuracy. Visualizations in Fig. 5 and Fig. 8 further demonstrate that SKNO excels in capturing sharp, local details, such as discontinuities at material interfaces and vorticity in low-viscosity fluids. For complete experimental settings, case studies, and complete results, see Appendix C.
[…]
@@ Results
Motivated by the assertion that the preparation and measurement of quantum states influence the outcomes in quantum mechanics, we explore different lifting and recovering operators to test their impact on the output solution. Here, the lifted $d+1$ dimensional function describes a quantum state as (Jin et al. (2023)). First, we fix $\mathcal{P}$ with a linear layer and vary the integration weights $\chi(p)$ for $\mathcal{Q}$. Tab. 13 shows that all recovering designs proposed in Example 1 perform well on the heat equation, where the Mean and Linear can also be properly scaled since the symmetry of $w(p)$. Amazingly, some predefined recovering operators can still work on complex Darcy cases, where the complete $p$ evolution provides detailed local changes on $D_{x}$ after the evolution on complete $p$, see Fig.7. Combining results from Tab. 13 and Tab. 13 on 2D Darcy, we find that for functions with complex topology on $D_{x}$, a continuous non-linear transformation on $d+1$ dimensional signals, either after first lifting or before last recovering, can help it adjust the $d+1$ dimensional propagating structure. The best performance with MLP-Linear combination is consistent with the MLP-Linear implementation in Transolver (Wu et al. (2024)) for handling complex topology. What’s more, comparing the Transolver performance in Tab.7 with ours in Tab.13, we also conclude that our $d+1$ dimensional design captures the evolution better on 2D Darcy.
[…]
capposition=top,capposition=top
[…]
@@
Figure 5: Left: Prediction, error, and second derivative of ground truth (GT) visualization for 2D Darcy; Right: Model performance on the Burgers, grid size ranging from 128 to 8192.
5 Conclusion
In this work, we redefine a novel $d+1$-dimensional framework for neural operators, inspired by the Schrödingerisation method that reformulates classical $d$-dimensional PDEs within a one-dimension-higher domain. Building on this foundation, we proposed the Schrödingerised Kernel Neural Operator (SKNO), which employs an efficient $d+1$ dimensional kernel integral operator to flexibly capture the underlying evolution. Across ten benchmarks, SKNO achieves state-of-the-art accuracy with competitive computational efficiency. Furthermore, its demonstrated resolution invariance and comprehensive ablation studies confirm both the robustness and validity of the proposed design. Taken together, these results establish SKNO as a principled and effective extension of neural operator methods, while also motivating broader exploration of quantum-inspired operator learning frameworks.
@@ Appendix A Table of Notations
Appendix B Statement on LLM Usage
In this paper, we used a large language model (LLM) to provide suggestions on word choice, grammar, and logical consistency. While some sentences were lightly rephrased by the LLM, all scientific content remains original and authored by us.
Appendix C Experiment Settings and Results
@@
Denote the parameterized neural operator as $\mathcal{G}_{\theta}$, we aim to approximate the solution operator $\mathcal{G}$ in Equation 1 by optimizing the model parameters $\theta\in\Theta$ through relative $L_{2}$ loss below:
$\min_{\theta\in\Theta}\mathcal{L}(\theta)$ $:=\min_{\theta\in\Theta}\frac{1}{S}\sum_{i=1}^{S}\left[\frac{\lVert{\mathcal{G}}_{\theta}[a_{i}]-u_{i}\rVert_{2}}{\lVert u_{i}\rVert_{2}}\right]$ (16)
$=\min_{\theta\in\Theta}\frac{1}{S}\sum_{i=1}^{S}\left[\left(\frac{\sum_{j}{\mathcal{G}}_{\theta}[a_{i}](x_{j})-u_{i}(x_{j})}{\sum_{j}u_{i}(x_{j})}\right)^{\frac{1}{2}}\right],$ (17)
where $\Theta$, $S$, and $u_{i}=u_{i}(x)$ denote the parameter space, the number of function samples, and the $i$-th output function, respectively.
@@
    where $c=0.05$ as in (Shin et al. (2022)). We generate 1200 samples (1000 for the training set, and 100 for testing) in a denser grid with size $2^{10}=1024$, and down-sample them to size $2^{8}=256$. Our goal is to predict solution $u_{1}=u(x,t=1)$ from $u_{0}=u(x,t=0)$. Results are shown in Tab.6.
@@
    where $\beta$ is set with $1$. We use the data from (Takamoto et al. (2022)) with 2000 samples of $u_{0}=u(x,t=0)$. Our goal is to predict the next 10 time steps from the last 10 of the solution u(x, t). The number of training and testing samples is 1000 and 200, respectively. Results are shown in Tab.6
@@
    The 1D Burger’s Equation on the unit torus is formed as (Li et al. (2021)):
@@
    where $D_{u}=0.2097,D_{v}=0.105$ denote the diffusion coefficients respectively for $u,v$. And $F=0.03,k=0.062$ are the reaction parameters. We follow the data generation in (Kassaï Koupaï et al. (2024)) to make a training set with 200 samples of $u,v$ evolution and a testing set with 20 in a $32\times 32$ size grid with periodic boundary conditions. Our goal is to predict the next 10 time steps from the last 10 of $u,v$. See Tab.7 for results.
@@
    where $h=h(x,t)$ denotes the fluid depth, $\mathbf{u}=\mathbf{u}(x,t)$ is the 2D velocity field, $g_{r}$ is the (reduced) gravitational acceleration, $b=b(x)$ represents the bathymetry characterizing the bottom topography. For this benchmark, we employ the dataset from (Takamoto et al. (2022)), which is generated under Neumann boundary conditions. To assess resolution-invariant stability during training, models are trained on 500 samples with resolutions randomly drawn from $\{32,64,128\}$ and evaluated on 300 test samples at the highest resolution 128. The prediction task is to forecast the water depth over the next 10 time steps, given the previous 10. See Tab.18 for results.
@@
    where $f=1$. Our goal is to approximate the mapping operator from the coefficient distribution function a(x)a({x}) to the solution function u(x)u({x}) with zero Dirichlet boundary conditions (Li et al. (2021)). The results are based on models trained on 1000 training samples, which are down-sampled to a size of ($85\times 85$), and tested on 100 samples. Results are listed in Tab.18.
@@
    Here $\nu=1e{-5}$ and $f=0.1(sin(2\pi(x_{1}+x_{2}))+cos(2\pi(x_{1}+x_{2})))$. Our goal is to predict the next 10 time steps from the last 10 ($d_{{a}}=d_{{u}}=10)$ of the solution w(x,t)w({x},t) with periodic boundary conditions on the 2D grid with size $64\times 64$ (Li et al. (2021)). All models in Table 7 are trained on 1000 training samples and tested on 200 test samples. Results are listed in Tab.7.
@@
    where $\textbf{v}=\textbf{v}(x,t)$, $\rho=\rho(x,t)$, and $p=p(x,t)$ denote the velocity vector field, fluid density (mass per unit volume), and pressure, respectively. We consider the challenging transonic flow benchmark from (Takamoto et al. (2022)), with shear and bulk viscosity coefficients $\eta=\zeta=1\times 10^{-8}$, and internal energy density $\epsilon=\tfrac{3p}{2}$. The viscous stress tensor $\sigma^{\prime}=\sigma^{\prime}(\textbf{v};\eta,\zeta)$ is determined by the velocity field under the Newtonian fluid assumption. Our objective is to predict the velocity, density, and pressure fields three time steps ahead, given the previous three steps ($d_{a}=d_{u}=3\times 5=15$). The simulations are performed on a downsampled 3D grid of size $32\times 32\times 32$ with periodic boundary conditions. The initial conditions consist of turbulent velocity fields with uniform mass density and pressure. All benchmark models are trained on 500 samples and evaluated on 100 test samples. See Tab.18 for results.

### v3

@@ 3.1 Model Architecture and Implementation
An overview of the SKNO architecture is shown in Fig. 3. Except for the lifting and measuring modules, we implement the SKNO by constructing signal propagators with the kernel integration in Equ.8, and with the aid of residuals. Each of them propagates the $d+1$ dimensional signal $\bm{v}_{l}$ with different structural biases on global or local regions on $D_{x}$.
@@
In the Schrödingerisation process (Jin et al., 2023) like Example 1 and 2, different differential operators with respect to $x$ (e.g. $\partial_{x}$, $\partial_{xx}$, …) can be represented by a complex diagonal matrix in Fourier domain of $x$ according to the properties of Fourier Transform.
Based on that, we leverage the implementation of spectral convolution operator in (Li et al., 2021), which uses a learnable complex vector equivalent to a diagonal matrix in the Fourier domain of $D_{x}$, to parameterize the evolving operator related to $d+1$ dimensional PDEs. A typical example is shown in 6. In practice, the complex vector is truncated for working on different grid sizes:
$\mathrm{SpectralConvOp}(\bm{v}_{l})=F_{x}^{-1}(\mathrm{trun\_diag}(F_{x}(\bm{v}_{l}))),$ (11)
where $F_{x}$ and $F_{x}^{-1}$ denote DFT and Inverse DFT (accelerated by Fast Fourier Transform (FFT) in practice). The $\mathrm{trun\_diag}()$ represents the element-wise multiplication between a learnable truncated complex vector and signal slices of $\bm{v}_{l}$ on the transformed grid $\xi$.
@@
Since the truncation in Sec.3.1 prefers to propagate signals in a relatively global scope, we compensate for possible local information lost by propagating the finite difference derivative of $x$ on the local region. Details are provided in Appendix B.
@@
With above two implementations of $d+1$ dimensional kernel integral operators in Equ.8, we specified the linear block $\mathcal{L}$ in Equ.2.2 with two parts: kernel integral operator $\mathcal{K}_{l}$ and residual operator $\bar{\mathcal{W}}_{l}$:
@@
Figure 4: Evolution along the auxiliary dimension and corresponding neural operator implementations. Left: The target operator $\mathcal{L}^{\prime}$ governs the evolution along pp. Right: Whereas prior methods rely on discretized matrices at each $p$ location, our model captures this evolution using operator modules acting on dual spaces.
@@ 3.2 Evolution along the Auxiliary Dimension
where $x_{{i}}=(x_{i_{1}},...,x_{i_{d}}),\ {i}=(i_{1},...,i_{d})$. Note that the same conclusion also works for $\xi_{i}$. And $K$ represents the highest order of the differential operator with respect to $p$.
After discretizing time $t$ across deep layers (Chen et al., 2018), Equ. 13 is implemented using a forward finite-difference (Euler) update. The step size is set to $\Delta t=1$ for notational simplicity, and can be treated as learnable or tuned in practice.
@@
which is exactly our linear block in Equ. 12. For discrete signals along $p$, we implement the right-hand side of Equ. 14 as ${F}_{p}^{-1}\tilde{A}{F}_{p}+\bm{b}$, where $\tilde{A}$ denotes a learnable matrix in the Fourier domain over $p$, and $\bm{b}$ corresponds to the discretized $b(p)$. As illustrated in Fig. 4, while other methods vary the discretization resolution of (p) to capture the evolving dynamics in Equ. 13, our model explicitly incorporates both the weighted differential operator represented in the Fourier domain over $p$ (SKNO, right) and the bias term (SKNO, left), thereby aligning with the general linear PDE form in Equ. 13. Extended implementations along $p$ are provided in Appendix B.
By iteratively updating the $d+1$ dimensional signal using the combination of the operator designs in Sec.3.1 and Sec.3.2, our SKNO model aligns itself well with $d+1$ dimensional PDEs, which effectively offers improved scalability and adaptability over previous methods.
@@ 4.1 General Setting
We evaluate the proposed model on more than ten diverse benchmarks spanning a wide spectrum of PDE dynamics. To highlight the effectiveness of our evolving kernel integral design, we first compare different kernel parameterizations on the PDE operators introduced in Sec. 2.1, including the 1D heat equation (Shin et al., 2022) and the 1D advection equation (Takamoto et al., 2022). We then evaluate full models on a broader set of PDE systems, conducting extensive experiments on the 1D Burgers equation; 2D Darcy flow with discontinuous coefficients; 2D incompressible Navier–Stokes at low viscosity (1e51e{-5}(Li et al., 2021; Wu et al., 2024a); 2D Gray–Scott Reaction–Diffusion (Kassaï Koupaï et al., 2024); and the highly unstable 3D Rayleigh–Taylor Instability (Ohana et al., 2024). Additional benchmarks and detailed results are provided in Appendix B.
@@
We compare our model with various $d$ dimensional neural implementations, including DeepONet (Lu et al., 2021), Fourier Neural Operator (FNO) (Li et al., 2021), Convolutional Neural Operator (CNO) (Raonic et al., 2023a), and transformer-based model Transolver (Wu et al., 2024a). All models are trained under their officially recommended optimal configurations, following the settings reported in their respective papers and official codes. For benchmarks such as 1D Burgers, 2D Darcy Flow, and 2D Gray–Scott, we train each model for 500 epochs with no physical prior regularization, using only relative $L_{2}$ loss. For the incompressible 2D Navier–Stokes, we ensure fairness by matching comparable training time scales, accounting for differences in training speed between transformer-based and other operator architectures. All experiments are run on a single Nvidia GeForce RTX 4090 GPU. Complete configurations are provided in Tab.7.
Table 1: Different $\mathcal{L}$ parameterizations’ performance on learning operators governed by 1D Heat and Advection Equations.
Linear Blocks Error
1D Heat Equation
FNO (1 mode, 4 grid $p$) 9.305e-1
Transolver (1 slice, 4 grid $p$ / 2 heads) 9.060e-1
Transolver (4 slices, 4 grid $p$ / 2 heads) 8.988e-1
Transolver (4 slices, 4 grid $p$ / 1 head) 8.987e-1
Ours (1 mode, 4 grid $p$) 2.637e-3
1D Advection Equation
Transolver (8 slices, 16 grid $p$ / 4 heads) 7.118e-1
Transolver (8 slices, 16 grid $p$ / 2 heads) 6.984e-1
Transolver (8 slices, 16 grid $p$ / 1 head) 6.828e-1
FNO (8 modes, 16 grid $p$) 6.012e-2
Ours (8 modes, 16 grid $p$) 1.979e-2
@@
4.2 Main Results and Analysis
Results
SKNO consistently outperforms existing methods, achieving the lowest relative $L_{2}$ error with competitive training cost. Tab. 1 summarizes the performance of different $\mathcal{L}$ parameterizations across representative models: the Fourier integral operator in FNO, the physics-attention block in Transolver, and SKNO’s spectral convolution operator acting on $d+1$ dimensional functions. All modules are evaluated with their original lifting and recovering components and without additional positional encoding (see Appendix B).
For the heat equation experiment, the solution at $t=1$ is predicted following Example 1, using the same truncation settings as in practice (i.e., truncated modes or slices, both of which parameterize an invertible aggregation on $D_{x}$). For the advection equation, these kernels are tested in an autoregressive setting to validate performance on sequential prediction with concatenated input quantities. Specifically, the reported error corresponds to predicting the next 10 time steps autoregressively given the previous 10.
While other modules are constrained by their evolution design along $p$, SKNO removes this bottleneck and stably captures the $d+1$ dimensional evolution. Across the benchmarks in Tab. 2, Tab. 13, and Tab. 15, SKNO achieves state-of-the-art (SOTA) performance, with training time comparable to FNO, which is widely recognized for balancing speed and accuracy. Visualizations in Fig. 9 and Fig. 10 further demonstrate that SKNO excels in capturing sharp, local details, such as discontinuities at material interfaces and vorticity in low-viscosity fluid flows. For complete experimental settings, case studies, and complete results, see Appendix B.
@@ Results
Table 3: Ablation study for SKNO on 2D Darcy. The Linear Residual is $\bar{\mathcal{W}}_{l}$ in Linear Block $\mathcal{L}$, while the non-linear one is out of $\mathcal{L}$.
Configuration Error
w.o. Double Res. 1.363e-2
w.o. Linear Res. 8.284e-3
w.o. Non-linear Res. 7.621e-3
w.o. $\tilde{A}$ along $p$ dim 8.804e-3
w.o. $\bm{b}$ along $p$ dim 6.827e-3
w.o. Global Propagators 1.584e-1
w.o. Local Propagator 5.715e-3
Baseline (SKNO) 5.555e-3
@@
We explore a broader set of lifting and recovering operators we have discussed to test their impact on the output solution. Specifically, we fix $\mathcal{P}$ with a linear layer and vary the integration weights $\chi(p)$ for $\mathcal{Q}$. Tab. 5 shows that all recovering designs proposed in Example 1 perform well on the heat equation, where the Mean and Linear can also be properly scaled since the symmetry of $w(p)$. Amazingly, some predefined recovering operators can still work on complex Darcy cases, where the complete evolution provides detailed local changes reflecting on $D_{x}$, see Fig.8. Combining results from Tab. 5 and Tab. 5 on 2D Darcy, we find that for functions with complex topology on $D_{x}$, a continuous non-linear transformation on $d+1$ dimensional signals, either after first lifting or before last recovering, can help it adjust the $d+1$ dimensional propagating structure. The best performance with MLP-Linear combination is consistent with the MLP-Linear implementation in Transolver (Wu et al., 2024a) for handling complex topology.
Table 4: The error on 1D Heat and 2D Darcy with different $\mathcal{Q}$ implementations.
$\mathcal{Q}$ Delta Step Mean Linear MLP MLP (Dropout)
1D Heat Equation
Error 2.538e-3 2.538e-3 2.538e-3 2.538e-3 2.637e-3 7.928e-3
2D Darcy Flow
Error 5.752e-3 1.402e-2 5.717e-3 5.741e-3 5.555e-3 5.678e-3
Table 5: The error on 2D Darcy with different $\mathcal{P}$, $\mathcal{Q}$ combinations.
$\mathcal{P}$ $\mathcal{Q}$ Linear MLP
Constant 8.738e-3 8.791e-3
Linear 5.741e-3 5.555e-3
MLP 5.952e-3 6.614e-3
MLP (Dropout) 5.451e-3 5.845e-3
@@
Other Experiments
More experiment results and visualizations can be found in the Appendix, mainly including (i) Zero-shot Generalization to Unseen Temporal Regimes: We have shown the SOTA performance of SKNO on 2D incompressible Navier-Stokes in Fig.15. (ii) Other Operators along $p$: We show that other operator choices along $p$ like physics-attention (Wu et al., 2024a) also benefit from adopting the $d+1$ dimensional operator design, see Tab.11.
Figure 5: Model performance on the Burgers equation with grid size ranging from 128 to 8192.
5 Conclusion, Limitations and Future work
In this work, we redefine neural operators within a $d+1$ dimensional framework. Building on this formulation, we propose the Schrödingerised Kernel Neural Operator (SKNO), which employs an efficient parameterization of the $d+1$ dimensional kernel integral operator. Across more than ten benchmarks, SKNO achieves state-of-the-art accuracy with competitive computational efficiency. Moreover, its demonstrated resolution invariance, zero-shot generalization to unseen temporal regimes, and comprehensive ablation studies jointly support the robustness and validity of the proposed $d+1$ dimensional designs. Taken together, these results not only establish $d+1$ dimensional neural operators such as SKNO as a simple yet highly effective extension of prior neural operators, but also motivate broader exploration of operator design and learning over the full $d+1$ problem domain.
An important open problem is to develop quantitative criteria for comparing neural operators at their optimizer-attained local minima. Because training navigates a high-dimensional error landscape over a large set of kernel parameters, both the optimization trajectory and the attained minimum can vary substantially across architectures. While existing models (e.g., (Li et al., 2021; Hao et al., 2023; Wu et al., 2024a)) and related theoretical analyses (e.g., (Kovachki et al., 2021; Lanthaler et al., 2025; Furuya et al., 2025)) provide approximation-oriented results, they offer limited insight into the properties of the minima selected by practical optimization. As future work, we will focus on comparing different aggregation designs and the associated propagation parameters across layers, and quantify how these choices affect neighborhood-averaged training and test errors around local minima.
6 Ethics and Reproducibility Statements
This work does not involve human subjects, personal data, or other sensitive information. All experiments were conducted on publicly available or synthetically generated datasets. To ensure reproducibility, we provide detailed descriptions of the experimental setup, including model architectures, training procedures, and evaluation protocols. The code and datasets used in this study will be made publicly accessible upon publication. For readers, the code implementations and related configurations are included in the supplementary materials.
@@ Appendix A Table of Notations
Others
$\hat{(\cdot)}$ The original function/signal represented in Fourier domain of $x$
$\tilde{(\cdot)}$ The original function/signal represented in Fourier domain of $p$
$\widehat{(\cdot)}$ The original operator/matrix represented in Fourier domain of $x$
$\widetilde{(\cdot)}$ The original operator/matrix represented in Fourier domain of $p$
$N_{x}$ The number of grid indices on $D_{x}$
$N_{p}$ The number of grid indices on $D_{p}$
$k_{x}$ The number of grid indices on the transformed domain after aggregation on $D_{x}$
$k_{p}$ The number of grid indices on the transformed domain after aggregation on $D_{x}$
Appendix B Model, Experiment Settings and Results
Differential Operator
We compensate for possible local information lost by propagating the finite difference derivative of $x$ on the local region:
where $D_{loc}\subset D_{x}$ denotes the local region around the query position $x$. In this work, the implementation of Equ.15 is based on the convolution layer in (Liu-Schiaffini et al., 2024). Related ablation results see Tab.3 and Tab.16.
@@ Appendix B Model, Experiment Settings and Results
Denote the parameterized neural operator as $\mathcal{G}_{\theta}$, we aim to approximate the solution operator $\mathcal{G}$ in Equation 2.2 by optimizing the model parameters $\theta\in\Theta$ through relative $L_{2}$ loss below:
$\min_{\theta\in\Theta}\mathcal{L}(\theta)$ $:=\min_{\theta\in\Theta}\frac{1}{S}\sum_{s=1}^{S}\left[\frac{\lVert{\mathcal{G}}_{\theta}[a_{s}]-u_{s}\rVert_{2}}{\lVert u_{s}\rVert_{2}}\right]$ (16)
$=\min_{\theta\in\Theta}\frac{1}{S}\sum_{s=1}^{S}\left[\left(\frac{\sum_{j}({\mathcal{G}}_{\theta}[a_{s}](x_{j})-u_{s}(x_{j}))^{2}}{\sum_{j}(u_{s}(x_{j}))^{2}}\right)^{\frac{1}{2}}\right],$ (17)
where $\Theta$, $S$, and $u_{s}=u_{s}(x)$ denote the parameter space, the number of function samples, and the $i$-th output function, respectively.
@@
    where $c=0.05$ as in (Shin et al., 2022). We generate 1200 samples (1000 for the training set, and 100 for testing) in a denser grid with size $2^{10}=1024$, and down-sample them to size $2^{8}=256$. Our goal is to predict solution $u_{1}=u(x,t=1)$ from $u_{0}=u(x,t=0)$. Results are shown in Tab.1.
@@
    where $\beta$ is set with $1$. We use the data from (Takamoto et al., 2022) with 2000 samples of $u_{0}=u(x,t=0)$. Our goal is to predict the next 10 time steps from the last 10 of the solution u(x, t). The number of training and testing samples is 1000 and 200, respectively. Results are shown in Tab.1
@@
    The 1D Burger’s Equation on the unit torus is formed as (Li et al., 2021):
@@
    where $D_{u}=0.2097,D_{v}=0.105$ denote the diffusion coefficients respectively for $u,v$. And $F=0.03,k=0.062$ are the reaction parameters. We follow the data generation in (Kassaï Koupaï et al., 2024) to make a training set with 200 samples of $u,v$ evolution and a testing set with 20 in a $32\times 32$ size grid with periodic boundary conditions. Our goal is to predict the next 10 time steps from the last 10 of $u,v$. See Tab.2 for results.
@@
    where $h=h(x,t)$ denotes the fluid depth, $\mathbf{u}=\mathbf{u}(x,t)$ is the 2D velocity field, $g_{r}$ is the (reduced) gravitational acceleration, $b=b(x)$ represents the bathymetry characterizing the bottom topography. For this benchmark, we employ the dataset from (Takamoto et al., 2022), which is generated under Neumann boundary conditions. To assess resolution-invariant stability during training, models are trained on 500 samples with resolutions randomly drawn from $\{32,64,128\}$ and evaluated on 300 test samples at the highest resolution 128. The prediction task is to forecast the water depth over the next 10 time steps, given the previous 10. See Tab.13 for results.
@@
    where $f=1$. Our goal is to approximate the mapping operator from the coefficient distribution function a(x)a({x}) to the solution function u(x)u({x}) with zero Dirichlet boundary conditions (Li et al., 2021). The results are based on models trained on 1000 training samples, which are down-sampled to a size of ($85\times 85$), and tested on 100 samples. Results are listed in Tab.13.
@@
    Here $\nu=1e{-5}$ and $f=0.1(sin(2\pi(x_{1}+x_{2}))+cos(2\pi(x_{1}+x_{2})))$. Our goal is to predict the next 10 time steps from the last 10 ($d_{{a}}=d_{{u}}=10)$ of the solution w(x,t)w({x},t) with periodic boundary conditions on the 2D grid with size $64\times 64$ (Li et al., 2021). All models in Table 2 are trained on 1000 training samples and tested on 200 test samples. Results are listed in Tab.2.
@@
    where $\textbf{v}=\textbf{v}(x,t)$, $\rho=\rho(x,t)$, and $p=p(x,t)$ denote the velocity vector field, fluid density (mass per unit volume), and pressure, respectively. We consider the challenging transonic flow benchmark from (Takamoto et al., 2022), with shear and bulk viscosity coefficients $\eta=\zeta=1\times 10^{-8}$, and internal energy density $\epsilon=\tfrac{3p}{2}$. The viscous stress tensor $\sigma^{\prime}=\sigma^{\prime}(\textbf{v};\eta,\zeta)$ is determined by the velocity field under the Newtonian fluid assumption. Our objective is to predict the velocity, density, and pressure fields three time steps ahead, given the previous three steps ($d_{a}=d_{u}=3\times 5=15$). The simulations are performed on a downsampled 3D grid of size $32\times 32\times 32$ with periodic boundary conditions. The initial conditions consist of turbulent velocity fields with uniform mass density and pressure. All benchmark models are trained on 500 samples and evaluated on 100 test samples. See Tab.13 for results.